A toy rocket (m =0.80 kg), initially at rest, is launched vertically upward such that over a time of 7.0 s the toy rocket travels a vertical displacement of 18 m. If the force of gravity on the toy rocket is 7.84 N, calculate the upward applied force of the engines. (Ignore air resistance)


1
Expert's answer
2021-01-16T17:24:17-0500

"F-F_g=ma\to F=F_g+m\cdot\frac{2s}{t^2}\to"


"F=F_g+m\cdot\frac{2s}{t^2}=7.84+0.8\cdot\frac{2\cdot18}{7}\approx12(N)"